Marula enterprises has developed a new local brew called Stiff Neck. The market demand for this product is given as follows: Q=240-4P. If stick Neck is priced at 40$, what is the point elasticity of demand?
1
Expert's answer
2020-04-08T09:35:00-0400

"Ep=Q'(P)×\frac{P1}{Q1};"


"Q'(P)=-4"

"Q1" "=240-4\u00d740;"

"Q1=240-160;"

"Q1=80;"

"P1=40;"

at a price of $ 40, demand is $ 80.

"Ep=-4\u00d740\/80=-2" %;